# Dynamic Obstacle Avoidance of Unmanned Surface Vehicles in Maritime Environments Using Gaussian Processes Based Motion Planning

### v1

@@ 4 Factor graph
In the proposed D-GPMP2, factor graph is used as a optimisation tool to deal with the MAP estimation and a new structure of it is designed to calculate the optimised trajectory to avoid static and dynamic obstacles simultaneously. More specifically, this new structure includes three categories of factors [1, 19]:
@@
    Prior factors such as 1) the regular prior factors and 2) the GP prior factors.
@@
    Static obstacle likelihood factors such as 1) the regular static obstacle likelihood factors and 2) the interpolated static obstacle likelihood factors.
@@
    Dynamic obstacle likelihood factors such as 1) the regular dynamic obstacle likelihood factors and 2) the interpolated dynamic obstacle likelihood factors.
@@ 5 Simulations and discussions
This section demonstrates the performance of the proposed motion planning algorithm on dynamic obstacle avoidance under different situations.
@@
5.1 Simulation details
In general, the proposed motion planning algorithm has been tested using five different categories of benchmark simulations:
  • 1.
    The first benchmark simulation validated the performance of the proposed motion planning algorithm when different weights of dynamic obstacle likelihood were chosen.
  • 2.
    The second benchmark simulation validated the performance of the proposed motion planning algorithm when the velocity of the dynamic obstacle was adjusted. More specifically, two quantitative sub-simulations were conducted with only the speed or direction of the dynamic obstacle adjusted.
  • 3.
    The third benchmark simulation validated the performance of the proposed motion planning algorithm when the area of the dynamic obstacle was adjusted.
  • 4.
    The fourth benchmark simulation validated the performance of the proposed motion planning algorithm in different multiple dynamic obstacles situations.
  • 5.
    The fifth benchmark simulation validated the performance of the proposed motion planning algorithm when both static and dynamic obstacles required to be avoided in complex maritime environments.
In all the benchmark simulations, the proposed motion planning algorithm was initialised with a CV-SL trajectory based on the system dynamics model described in Section 5.2. Table 3 demonstrates the specification of the used hardware platform for running the benchmark simulations, while Fig. 10 demonstrates the maritime map used in the fifth benchmark simulation. Meanwhile, the obstacle angle in the following benchmark simulations is the target USV course.
@@ 5.3 Simulation with different weights of dynamic obstacle likelihood
5.4 Simulation with a dynamic obstacle with different velocities
@@ 9 Safe radius 9 [m]
5.4.1 Simulation with a dynamic obstacle with different speeds
@@ 8 Safe radius 9 [m]
5.4.2 Simulation with a dynamic obstacle with different directions
@@ 5.6 Simulation with multiple dynamic obstacles
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [20, 20]
3 Goal position [80, 80]
4 Obstacle speed 3 [m/s]
5 Obstacle length 1.5 [m]
6 Obstacle width 1 [m]
7 Safe radius 2.5 [m]
8 Dynamic obstacle likelihood weight 0.005
@@ 12 Safe distance 20 [m]
5.7 Simulation with static and dynamic obstacles.
@@
6 Implementation in ROS
6.1 Implementation details
A widely-used catamaran, the WAM-V 20 USV, is provided by the high-fidelity maritime environment and this environment is provided by [27]. Fig. 17 provides an intuitive overview of the WAM-V 20 USV above and below sea level in the ROS Gazebo virtual world. Two WAM-V 20 USVs were built inside it for a dynamic obstacle avoidance mission. As demonstrated in Fig. 18, the proposed D-GPMP2 algorithm was run on MATLAB and connected with the WAM-V 20 USVs in ROS Gazebo through ROS node.
Based on the aforementioned benchmark simulations, the proposed D-GPMP2 algorithm is able to generate an optimised path to avoid different types of dynamic obstacles. In reality or high-fidelity simulation environments, following a generated path to avoid a dynamic obstacle does not happen in a flash and takes time. As a result, we increase the input speed of a dynamic obstacle by adding a scaling factor. Despite that following a generated path consumes time, this would help guide our WAM-V 20 USV to avoid the dynamic obstacle.
@@ 6.2 Dynamic obstacle avoidance mission
During the dynamic obstacle avoidance mission, our WAM-V 20 USV transports a package from a start position to a goal position, while another WAM-V 20 USV moves towards our WAM-V 20 USV with a constant velocity. We use the proposed D-GPMP2 algorithm to generate an optimised path based on the real-time status information of the dynamic obstacle when transporting the package. By following the optimised path generated by the proposed D-GPMP2 algorithm, our WAM-V 20 USV successfully avoids the dynamic obstacle during the transportation period. In this mission, we assume the USVs have ”perfect” autopilots, meaning they follow the designated paths perfectly.
To provide a more intuitive understanding of the above process, the storyboards of the dynamic obstacle avoidance mission from the third-person perspectives is demonstrated in Fig. 19.
@@ 7 Conclusion
The work in this article successfully extended the traditional GP-based motion planning algorithm into complex and dynamic maritime environments. To achieve this purpose, a novel motion planning algorithm is proposed in this article, named as the D-GPMP2. By introducing 1) dynamic obstacle area modelling and 2) dynamic obstacle velocity modelling, we were able to use the proposed D-GPMP2 algorithm to generate safe, smooth and relatively short trajectories for USVs to avoid both static and dynamic obstacles in maritime environments in an efficient manner. Moreover, a variety of simulation environments across different platforms have been used to test the practicality and functionality of the proposed D-GPMP2 algorithm.
7.1 Future research
In terms of future research, proposed areas of focus are summarised as below:
  1. 1.
    Developing the proposed D-GPMP2 algorithm further to make it usable for AUVs in 3-D environments.
  2. 2.
    Testing the proposed D-GPMP2 algorithm on a real physical USV to further validate its practicality.

### v2

@@ 5 Factor graph
In the proposed D-GPMP2 and D-GPMP2-COLREGs algorithms, factor graph is used as a optimisation tool to deal with the MAP estimation and a new structure of it is designed to calculate the optimised trajectory to avoid static and dynamic obstacles simultaneously. More specifically, this new structure includes three categories of factors [1, 19]:
@@
    Prior factors such as 1) the regular prior factors [$f^{p}_{0}(\theta_{0})$ and $f^{p}_{N}(\theta_{N})$] and 2) the GP prior factors [$f^{gp}_{i}(\theta_{i},\theta_{i+1})$].
@@
    Static obstacle likelihood factors such as 1) the regular static obstacle likelihood factors [$f^{st\_obs}_{i}(\theta_{i})$] and 2) the interpolated static obstacle likelihood factors [$f^{st\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$].
@@
    Dynamic obstacle likelihood factors such as 1) the regular dynamic obstacle likelihood factors [$f^{dy\_obs}_{i}(\theta_{i})$] and 2) the interpolated dynamic obstacle likelihood factors [$f^{dy\_obs}_{\tau_{j}}(\theta_{i},\theta_{i+1})$].
@@ 6 Simulations and discussions
This section demonstrates the performance of the proposed motion planning algorithms on dynamic obstacle avoidance under different situations in a series of qualitative and quantitative simulations.
6.1 Simulation details
In general, the proposed motion planning algorithms have been tested using seven different categories of benchmark simulations:
  • 1.
    The first benchmark simulation validated the performance of the proposed D-GPMP2 algorithm when different weights of dynamic obstacle likelihood were chosen.
  • 2.
    The second benchmark simulation validated the performance of the proposed D-GPMP2 algorithm when the velocity of the dynamic obstacle was adjusted. More specifically, two quantitative sub-simulations were conducted with only the speed or direction of the dynamic obstacle adjusted.
  • 3.
    The third benchmark simulation validated the performance of the proposed D-GPMP2 algorithm when the area of the dynamic obstacle was adjusted.
  • 4.
    The fourth benchmark simulation validated the performance of the proposed D-GPMP2 algorithm in different multiple dynamic obstacles situations.
  • 5.
    The fifth benchmark simulation validated the performance of the proposed D-GPMP2 algorithm when both static and dynamic obstacles required to be avoided in complex maritime environments.
  • 6.
    The sixth benchmark simulation compared the performance of the proposed D-GPMP2 algorithm with an existing method (DWA) in complex maritime environments.
  • 7.
    The seventh benchmark simulation demonstrated the performance of the proposed D-GPMP2-COLREGs algorithm in overtaking, head-on and crossing scenarios.
@@ 6.1 Simulation details
In all the benchmark simulations, the proposed motion planning algorithms were initialised with a CV-SL trajectory based on the system dynamics model described in Section 6.2. Table 4 demonstrates the specification of the used hardware platform for running the benchmark simulations, while Fig. 11 demonstrates the maritime map used in the fifth benchmark simulation. Meanwhile, the obstacle angle in the following benchmark simulations is the target USV course. The parameters used in the following simulations were identified based on empirical measurements.
@@ 9 Safe radius 9 [m]
6.4 Simulation with a dynamic obstacle with different velocities
@@ 8 Safe radius 9 [m]
6.4.1 Simulation with a dynamic obstacle with different speeds
@@
6.4.2 Simulation with a dynamic obstacle with different directions
@@ 6 Obstacle speed 5 [m/s]
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [20, 20]
3 Goal position [80, 80]
4 Obstacle speed 3 [m/s]
5 Obstacle length 1.5 [m]
6 Obstacle width 1 [m]
7 Safe radius 2.5 [m]
8 Dynamic obstacle likelihood weight 0.005
@@ 12 Safe distance 20 [m]
6.8 Benchmark comparison with DWA
Figure 18: Demonstration of the simulation result of the benchmark comparison with DWA. (a) The trajectory generated by the proposed D-GPMP2 algorithm. (b) The trajectory generated by DWA. The start positions are represented as green circles, the goal positions are represented as red circles, the trajectory generated by the proposed D-GPMP2 algorithm is represented as black line, the trajectory generated by DWA is represented as blue line and the potential local trajectories of DWA are represented as green arrows. In the colour bar of sub-figure (a), a larger value indicates a higher risk and a smaller value indicates a lower risk.
Table 12: List of parameters of the proposed D-GPMP2 algorithm in benchmark comparison.
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [20, 50]
3 Goal position [80, 50]
4 Obstacle position [50, 50]
5 obstacle speed 5 [m/s]
6 Obstacle angle 270 [degree]
7 Obstacle length 6 [m]
8 Obstacle width 3 [m]
9 Safe radius 9 [m]
10 Likelihood weight 0.01
Table 13: List of parameters of DWA in benchmark comparison.
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [20, 50]
3 Goal position [80, 50]
4 Obstacle position [50, 50]
5 Safe radius 5 [m]
6 Sample interval 0.5 [s]
7 Prediction time 3 [s]
8 Linear velocity resolution 0.05 [m/s]
9 Angular velocity resolution 0.1 [rad/s]
Table 14: Summary of the simulation result in the benchmark comparison between the proposed D-GPMP2 and DWA.
Number Measurement D-GPMP2 DWA
1 Path length 64.4 [m] 67.8 [m]
2 Time cost 28.9 [ms] 68.2 [ms]
This subsection compares the performance of the proposed D-GPMP2 algorithm with an existing obstacle avoid strategy (DWA) in an environment with a dynamic obstacle.
Only a quantitative comparison with DWA is presented, as mainstream motion planning algorithms like A* and RRT* cannot meet real-time requirements for dynamic obstacle avoidance in many cases due to their computational inefficiency. Comprehensive benchmarking data highlighting the performance differences on computational speed between GP-based algorithms and mainstream motion planning algorithms are detailed in [1, 2].
This benchmark simulation compared the performance between the proposed D-GPMP2 algorithm and DWA in an obstacle avoidance scenario, where a moving obstacle exists between the start and goal positions. Table 12 demonstrates the parameters used in the proposed D-GPMP2 algorithm, while Table 13 demonstrates the parameters used in the DWA. To align with the dynamic obstacle modelling in D-GPMP2, the obstacle model in DWA was expanded along its movement direction. Fig. 18 demonstrates the simulation results of the proposed D-GPMP2 and DWA, while Table 14 compares both the generated path length and total time required for generating the path to avoid the moving obstacle. It is worth noting that D-GPMP2 can generate a shorter path with reduced computation time for dynamic obstacle avoidance compared with DWA.
In summary, the quantitative comparison with a mainstream obstacle avoidance strategy (DWA) demonstrates D-GPMP2’s superior performance in dynamic obstacle avoidance.
6.9 COLREGs-compliant benchmark simulation
This subsection provides a quantitative performance analysis of the proposed D-GPMP2-COLREGs algorithm for COLREGs-constrained dynamic obstacle avoidance in maritime environments.
The benchmark demonstrates that the D-GPMP2-COLREGs algorithm effectively handles dynamic obstacle avoidance in maritime scenarios while strictly complying with COLREGs regulations. Fig. 19 (a) demonstrates the result of the proposed D-GPMP2-COLREGs algorithm in an overtaking simulation scenario. Because COLREGs don’t strictly specify the overtaking side, our USV overtakes the encountered USV by passing on its starboard side. Fig. 19 (b) demonstrates the result of the proposed D-GPMP2-COLREGs algorithm in a head-on simulation scenario. In compliance with the COLREGs rule for head-on situations, which requires port-side passing, the proposed D-GPMP2-COLREGs algorithm generates a regulation-compliant trajectory. Fig. 19 (c) demonstrates the result of the proposed D-GPMP2-COLREGs algorithm in a crossing simulation scenario. As can be seen in this sub-figure, our USV passes behind the encountered vessel’s port side in compliance with COLREGs rules.
In summary, the proposed D-GPMP2-COLREGs algorithm can effectively avoid dynamic obstacles while following COLREGs rules.
Table 15: List of parameters of the proposed D-GPMP2-COLREGs algorithm in COLREGs-compliant benchmark simulation (overtaking).
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [50, 20]
3 Goal position [50, 80]
4 Obstacle position [50, 50]
5 obstacle speed 5 [m/s]
6 Obstacle angle 0 [degree]
7 Obstacle length 6 [m]
8 Obstacle width 3 [m]
9 Safe radius 9 [m]
10 Likelihood weight 0.01
Table 16: List of parameters of the proposed D-GPMP2-COLREGs algorithm in COLREGs-compliant benchmark simulation (head-on).
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [50, 20]
3 Goal position [50, 90]
4 Obstacle position [50, 70]
5 obstacle speed 3 [m/s]
6 Obstacle angle 180 [degree]
7 Obstacle length 6 [m]
8 Obstacle width 3 [m]
9 Safe radius 9 [m]
10 Likelihood weight 0.01
Table 17: List of parameters of the proposed D-GPMP2-COLREGs algorithm in COLREGs-compliant benchmark simulation (crossing).
Number Parameter name Parameter value
1 Map size 100 $\times$ 100 [m2]
2 Start position [50, 20]
3 Goal position [70, 80]
4 Obstacle position [65, 60]
5 obstacle speed 5 [m/s]
6 Obstacle angle 270 [degree]
7 Obstacle length 5 [m]
8 Obstacle width 2 [m]
9 Safe radius 7 [m]
10 Likelihood weight 0.01
Figure 19: Demonstration of the results of COLREGs-compliant benchmark simulation. (a) Overtaking result, (b) Head-on result and (c) Crossing result. In all sub-figures, our USV is positioned at the starting point and moving along the positive Y-axis direction. The green arrows indicate the moving directions of the dynamic obstacles.
7 Implementation in ROS
7.1 Implementation details
Compared with practical experiments in real world, the implementation in ROS has a lower cost and is more efficient. The implementation in ROS considers the physical properties such as gravity, buoyancy, hydrodynamic forces, inertia, deformation caused by collision as well as other necessary factors to achieve exceptional correlation with actual experimental results. A widely-used catamaran, the WAM-V 20 USV, is provided by the high-fidelity maritime environment and the environment is provided by [35]. Fig. 20 provides an intuitive overview of the WAM-V 20 USV above and below sea level in the ROS Gazebo virtual world. An additional WAM-V 20 USV was implemented in the simulation environment for the following dynamic obstacle avoidance mission. As demonstrated in Fig. 21, the proposed D-GPMP2 algorithm was run on MATLAB and connected with the WAM-V 20 USVs in ROS Gazebo through ROS node. In real-world scenarios and high-fidelity simulation environments, dynamic obstacle avoidance through path following requires non-negligible time cost and cannot be performed instantaneously.
@@ 7.2 Dynamic obstacle avoidance mission
During the dynamic obstacle avoidance mission, the primary USV completes a package delivery mission between designated positions while evading a secondary USV moving at constant velocity. We use the proposed D-GPMP2 algorithm to generate an optimised path based on the real-time status information of the dynamic obstacle when transporting the package. By following the optimised path generated by the proposed D-GPMP2 algorithm, the primary USV maintains successful dynamic obstacle avoidance throughout its entire transportation mission. During the dynamic obstacle avoidance mission, it is assumed the USVs owns ”perfect” autopilots, which indicates these USVs can follow the designated paths perfectly. To provide a more intuitive understanding of the above process, the storyboards of the dynamic obstacle avoidance mission from the third-person perspectives is demonstrated in Fig. 22.
@@ 8 Conclusion
The work in this article successfully extended the traditional GP-based motion planning algorithm into complex and dynamic maritime environments. To achieve this purpose, a novel motion planning algorithm is proposed in this article, named as the D-GPMP2. By introducing dynamic obstacle area modelling and dynamic obstacle velocity modelling, we were able to use the proposed D-GPMP2 algorithm to generate safe, smooth and relatively short trajectories for USVs to avoid both static and dynamic obstacles simultaneously in maritime environments in an efficient manner. Furthermore, D-GPMP2-COLREGs is proposed, which is an enhanced variant of D-GPMP2 that incorporates COLREGs compliance for dynamic obstacle avoidance in maritime environments. A variety of simulation environments across different platforms have been used to validate the practicality and functionality of the proposed algorithms.
More specifically, the contributions of this work can be summarised as follows:
  • 1.
    This article first presented an approach for constructing adaptive safe areas for moving USVs. This is accomplished through integration of sensor measurements of dimensional parameters into the multivariate Gaussian distribution.
  • 2.
    Then this article presented an approach for integrating the sensor measurements of the real-time motion state (speed, position and course) into the adaptive safe area. This is achieved through mathematical modelling innovations coupled with computer vision techniques based on the adaptive safe area.
  • 3.
    This article presented an approach for integrating COLREGs rules into the proposed motion planning algorithm, which further enhance its practicability in maritime environments.
  • 4.
    This article also presented a new factor graph structure for GP-based motion planning, capable of handling both static and dynamic obstacles simultaneously.
  • 5.
    Last, this article thoroughly tested and validated the proposed motion planning algorithms in both a series of self-constructed simulation environments and a high-fidelity maritime environment in ROS.
In terms of future research, some proposed areas of focus are summarised as: 1) extending the proposed D-GPMP2 algorithm to enable 3-D motion planning for AUVs and 2) validating the proposed D-GPMP2 algorithm through real-world USV experiments to further demonstrate its practicality.